OpenStudy (anonymous):

Suppose f(x)=−7x^8+30x^6−17. (a) Then f(−x)=? (Simplify) Transformations and shifts help

OpenStudy (anonymous):

x is raised to an even power in each case so f(x)=f(-x), it is an even function

OpenStudy (anonymous):

which means?

OpenStudy (anonymous):

reflected over the y axis

OpenStudy (anonymous):

not reflected, means it is symmetric with respect to y-axis

OpenStudy (anonymous):

so how do we simplify that problem

OpenStudy (anonymous):

simplify it?

OpenStudy (anonymous):

oh, you want the transformation....yes f(-x) is the graph of f(x) reflected across the y-axis, but the graph will look exactly the same considering that f(x) is already symmetric with respect to y-axis

OpenStudy (anonymous):

so what will my new equation be?

OpenStudy (anonymous):

is it just the same thing?

OpenStudy (anonymous):

f(x)=−7x^8+30x^6−17 f(-x)=-7(-x)^8+30(-x)^6-17 simplified means remove the negative sign where it isn't required. since (-x) raised to any even power is positive, you can remove both negatives f(-x)=-7x^8+30x^6-17

OpenStudy (anonymous):

bingo!

OpenStudy (anonymous):

−8x7+12x−12 f(-x) = ? would this be the same thing then?

OpenStudy (anonymous):

−8x7^+12x−12

OpenStudy (anonymous):

got it. Now i have a problem with this Write a formula (in terms of the independent variable w) for the transformation of k(w)=5^w given below. y=−k(w−3)= ???